A metal sphere has a charge of +8.0 µC.
What is the net charge after 6.0×10
13
electrons have been placed on it?
1
Expert's answer
2020-09-30T12:26:48-0400

One electron has charge "-1.6 \cdot 10^{-19}" C. Then placing "6.0\u00d710^{13}" electrons on the sphere will bring "- 9.6 \cdot 10^{-6} \, C". So, total charge becomes

"-9.6 \cdot 10^{-6} +8 \cdot 10^{-6} = -1.6 \cdot 10^{-6} \, C"

Answer: "-1.6 \cdot 10^{-6} \, C"